OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form

OpenStudy (asadkarim7):

take gradient = 3-2/ 6-2 1/4 y=mx+c 2= 1/4 ( 2) + c 2- 1/2 = c 3/2 = c y= 1/4 x + 3/2 4y = x + 6